# On prior smoothing with discrete spatial data in the context of disease mapping

### v1

@@ Acknowledgements
The authors thank Miguel Ángel Beltrán and Miguel Ángel Martínez-Beneito for useful conversation in the model fitting.
@@
The work was supported by Project PID2020-113125RB-I00/MCIN/AEI/10.13039/501100011033. Garazi Retegui is supported by PhD student scholarship from the Public University of Navarra together with Banco Santander (Ayudas Predoctorales Santander UPNA 2021-2022). Jaione Etxeberria and María Dolores Ugarte would like to acknowledge support from Project UNEDPAM/PI/PR24/05A.
Competing Interests
@@
Code Availability
The data and code supporting the findings of this study will be made openly available on GitHub.
@@ Appendix B Within prior simulation study
For the GP prior, as discussed in Section 2.1 of the main paper, the level of smoothing induced by the prior is influenced by the values of $\sigma^{2}$ and the range parameter $\psi$. Therefore, in this simulation study we vary the $\sigma^{2}$ and $\psi$ values. Specifically, we considered five values ranging from $10^{-4}$ to $0.25$ for $\sigma^{2}$ and three different options for $\psi$: $1$, $5$, and $9$. Table 7 shows the empirical smoothing criteria, MSS, RMSS and the maximum values, $maxMSS$ and $maxRMSS$, and the smoothing proportion (SP), together with the theoretical smoothing (TCV) for the different parameter values. Theoretical smoothing aligns with that determined through empirical metrics. Remind that theoretical measures demonstrate that as the $\sigma$ value increases, the expected smoothing decreases, while conversely, as the $\psi$ value increases, the smoothing increases. However, for the extreme values of $\sigma^{2}$, we appreciate different behavior for the $\psi$ parameter. Specifically, for $\sigma^{2}=10^{-4}$, changes in $\psi$ parameter slightly affect the amount of smoothing. Similar to the other priors analyzed, this suggests that the smoothing for the GP prior reaches a maximum value. Something similar happens for the empirical smoothing metrics with $\sigma^{2}=0.25$, while the theoretical measure consistently increases. Therefore, we observe that smoothing tends to an asymptote, as seen with the iCAR and LCAR priors in the main text. Across different numbers of areas, higher empirical values (MSS, RMSS, and maximum values) observed with $A=300$ result in lower smoothing and lower SP values compared to $A=47$. Additionally, for a TCV value of zero, the SP is lower for $A=300$ than for $A=47$ indicating that the maximum smoothing decreases as the number of areas increases. Moreover, a faster increase in TCV is observed with a greater number of areas, suggesting that smoothing approaches an asymptote more quickly as the spatial disaggregation increases.

### v2

@@ Acknowledgements
The authors thank Miguel Ángel Beltrán and Miguel Ángel Martínez-Beneito for useful conversation in the model fitting. The authors also thank the reviewers and the associate editor for their helpul comments.
@@
The work was supported by Project PID2020-113125RB-I00/MCIN/AEI/10.13039/501100011033, Project PID2024-155382OB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE and BIOSTATNET - PROYECTOS REDES DE INVESTIGACIÓN 2024 - RED2024-153680-T/MICIU/AEI/. Garazi Retegui is supported by PhD student scholarship from the Public University of Navarra together with Banco Santander (Ayudas Predoctorales Santander UPNA 2021-2022). Jaione Etxeberria and María Dolores Ugarte would like to acknowledge support from Project UNEDPAM/PI/PR24/05A.
Declaration of conflicting interests
@@
Data Availability
The data and code supporting the findings of this study are available at https://github.com/spatialstatisticsupna/Prior_Smoothing.
@@ Appendix B Within prior simulation study
For the GP prior, as discussed in Section 2.1 of the main paper, the level of smoothing induced by the prior is influenced by the values of $\sigma^{2}$ and the range parameter $\psi$. Therefore, in this simulation study we vary the $\sigma^{2}$ and $\psi$ values. Specifically, we considered five values ranging from $10^{-4}$ to 0.250.250.25 for $\sigma^{2}$ and three different options for $\psi$: 111, 555, and 999. Table 7 shows the empirical smoothing criteria, MSS, RMSS and the maximum values, $maxMSS$ and $maxRMSS$, and the smoothing proportion (SP), together with the theoretical smoothing (TCV) for the different parameter values. Theoretical smoothing aligns with that determined through empirical metrics. Remind that theoretical measures demonstrate that as the $\sigma$ value increases, the expected smoothing decreases, while conversely, as the $\psi$ value increases, the smoothing increases. However, for the extreme values of $\sigma^{2}$, we appreciate different behavior for the $\psi$ parameter. Specifically, for $\sigma^{2}=10^{-4}$, changes in $\psi$ parameter slightly affect the amount of smoothing. Similar to the other priors analyzed, this suggests that the smoothing for the GP prior reaches a maximum value. Something similar happens for the empirical smoothing metrics with $\sigma^{2}=0.25$, while the theoretical measure consistently increases. Therefore, we observe that smoothing tends to an asymptote, as seen with the iCAR and LCAR priors in the main text. Across different numbers of areas, higher empirical values (MSS, RMSS, and maximum values) observed with $A=300$ result in lower smoothing and lower SP values compared to $A=47$. Additionally, for a TCV value of zero, the SP is lower for $A=300$ than for $A=47$ indicating that the maximum smoothing decreases as the number of areas increases. Moreover, a faster increase in TCV is observed with a greater number of areas, suggesting that smoothing approaches an asymptote more quickly as the spatial disaggregation increases.